Earth presently has 1.4 billion hectares of arable land, and each hectare produces between 1 and 2 tons of grain annually. What is the amount of grain (in tons/yr) produced annually? 2.1e9 Correct: Your answer is correct. tons/yr If every human being became vegetarian (as feeding livestock and consuming the products of livestock is far less efficient), then 200 kg of grain would be sufficient to feed each human being on the planet. Based on this data, what is the number of people Earth can support? (Use your value for the amount of grain produced annually.) 9.525e10 Incorrect: Your answer is incorrect. How did you obtain your estimate? How many kilograms per person does the problem statement indicate? people
Answer #1

1 tonn =1000 kg

Grain produced per year = 2.1×10^12 kg

1 person needs =200kg

Hence the no of people = 2.1×10^12 /200 = 105×10^7 people.